OpenStudy (yanni):

The length of a rectangle is 2ft longer than its width. If the perimeter of the rectangle is 40ft , find its area.

OpenStudy (anonymous):

\[2w + 2(w+2) = 40\]

OpenStudy (yanni):

2w+2w+4=40?

OpenStudy (anonymous):

this should help you compute w and l

OpenStudy (anonymous):

\[l(length)= w+2\]

OpenStudy (yanni):

i dont get it

OpenStudy (anonymous):

\[A _{rectangle}= l * w\] \[w=width,l=length\] \[length=width + 2ft\] \[width + width +(width+2) +(width+2) = 40ft\]

OpenStudy (anonymous):

solve for w and you will get the value of the width. add 2 ft and you will get the value of the length. compute area using these values.